# When large language models meet evolutionary algorithms

### v2

@@ 2 Evolutionary fine-tuning in black-box scenarios
Evolutionary model tuning targets parameter space, while evolutionary prompt tuning and self-tuning focus on language (search) space. Compared to model training involving multiple gradient descents, evolutionary black-box tuning is highly cost-effective. Current research focuses on model evolution within the language space. In open environments, complex tasks may require self-coevolution in language and parameter spaces, posing challenges like resource management, catastrophic forgetting, fitness assessment, and security issues. Efficient resource management strategies help save costs in continuous evolution. Finding a balance between new and old knowledge mitigates catastrophic forgetting. Designing collaborative evaluation strategies for language and parameter spaces tailored to specific tasks is essential. Self-evolving systems urgently require a robust security assessment mechanism to avert ethical challenges. In addition, integrating text, images, audio, and video is increasingly crucial [83]. Evolutionary multimodal fusion techniques [82, 65] offer a promising path to unifying diverse information, thereby expanding the applicability and versatility of evolutionary fine-tuning.
@@ 4 Conclusion
LLMs and EAs have spurred innovation across interdisciplinary domains. The synergistic integration of EAs with LLMs is poised to realize the evolutionary learning machines proposed by Turing [39]. This paper demonstrates the parallels between LLMs and EAs from five aspects, initially indicating that analogies can potentially spark a new artificial intelligence paradigm integrating LLM’s learning abilities with EA’s exploratory capabilities. Recently, LLMs have shown promise in utilizing principles of evolution [31, 4, 27, 85, 66, 67, 68, 69, 70]. The exponential growth in computing power enables large models combined with evolutionary mechanisms to perform reasoning in complex environments. Analogies offer an early explanation for interconnected studies.

### v3

@@ 2 Evolutionary fine-tuning in black-box scenarios
Evolutionary model tuning targets parameter space, while evolutionary prompt tuning and self-tuning focus on language (search) space. Compared to model training involving multiple gradient descents, evolutionary black-box tuning is highly cost-effective. Current research focuses on model evolution within the language space. In open environments, complex tasks may require self-coevolution in language and parameter spaces, posing challenges like resource management, catastrophic forgetting, fitness assessment, and security issues. Efficient resource management strategies help save costs in continuous evolution. Finding a balance between new and old knowledge mitigates catastrophic forgetting. Designing collaborative evaluation strategies for language and parameter spaces tailored to specific tasks is essential. As self-evolving systems continue to advance, the development of robust security assessment mechanisms may become critical to address potential ethical challenges. In addition, integrating text, images, audio, and video is increasingly crucial [94]. Evolutionary multimodal fusion techniques [93, 76] offer a promising path to unifying diverse information, thereby expanding the applicability and versatility of evolutionary fine-tuning.
@@ 4 Conclusion
LLMs and EAs have spurred innovation across interdisciplinary domains, with their synergistic integration holding the potential to realize the evolutionary learning machines envisioned by Turing [45]. This paper demonstrates the conceptual parallels between LLMs and EAs from five aspects, initially indicating that analogies can potentially spark a new artificial intelligence paradigm integrating LLM’s learning abilities with EA’s exploratory capabilities. Recently, LLMs have shown promise in utilizing principles of evolution [37, 9, 33, 96, 77, 78, 79, 80, 81]. The exponential growth in computing power enables large models combined with evolutionary mechanisms to perform reasoning in complex environments. Building on these developments, our work highlights promising directions for advancing current research and identifies critical challenges for future progress.